# A characterization of the local structure of two-dimensional sets with positive reach

### v1

@@ 1. Introduction
The following result is a consequence of statements claimed in [7, Remark 4.20] and easily follows from [23, Proposition 7.4].
@@
Recently, A. Lytchak proved an interesting result ([18, Theorem 1.2]) which is essentially an extension of Theorem F. Lytchak’s result deals with sets in $d$-dimensional Riemannian manifold; its (essentially equivalent) Euclidean version is the following.
@@
Note that Theorem L not only generalizes Theorem F (which corresponds to the case $\Phi^{-1}(a)\in{\rm int}\,K$), but also gives (for $k=1$) a local characterization of one-dimensional sets $A$ with positive reach in $\mathbb{R}^{d}$ which is equivalent to that which follows from [23, Corollary 8.9]: on a neighbourhood of $a\in A$, $A$ is either a singleton, or a simple $C^{1,1}$ arc.
If $A\subset\mathbb{R}^{2}$ has positive reach then its local description using semiconcave and semiconvex functions (without any further restrictions) follows easily from [23, Theorem 6.4]. This result, however, cannot be easily extended to higher dimensions.
The main result of the present article is a complete characterization of the local structure of two-dimensional sets $A$ with positive reach in $\mathbb{R}^{d}$ for $d\geq 3$. Note that Theorem L gives such a characterization at points $a\in A$ where $\operatorname{Tan}(A,a)$ has dimension $2$; the only nontrivial remaining case is when $\dim\operatorname{Tan}(A,a)=1$. This case, however, is much different from the case $\dim\operatorname{Tan}(A,a)=2$ - in particular, we cannot expect that $A$ will lie on a two-dimensional $C^{1,1}$ surface in some neighbourhood of $a$. This follows already from [23, Example 7.13], which we will describe now in a slightly generalized version.
@@
Our main result shows (rather surprisingly) that the above example is in a sense universal (up to a $C^{1,1}$-diffeomorphism) for the case when the tangent cone of $A$ at $a$ is a half-line. In our proof of the main result (Theorem 1.3 below), we need to include the “both-sided” version of the set $B$ above (where the tangent cone is the full line, see case (ii) in Definition 1.1), and to define properly the “multirotations” in $\mathbb{R}^{d}$. So, before stating the result, we need two definitions.
@@
Note that if $B$ is a B-set then, since it has positive reach, $T_{1}(B)$ is a compact subset of $\mathbb{R}=\operatorname{span}\{e_{1}\}$ by Lemma 3.3 (iii).
@@
Moreover, if $\Phi$, $B$ and $\rho$ are as above, then $\Phi(\rho(B))$ agrees on a neighbourhood of $a$ with some set $A\subset\mathbb{R}^{d}$ with positive reach satisfying $\dim A\leq 2$ and $\dim(\operatorname{Tan}(A,a))=1$.
@@
It is easy to see that Theorem 1.3 holds if we replace $\dim A\leq 2$” with $\dim A=2$” at both occurrences. The only difference is in the proof of the “moreover” part: While in the first case we set in our proof $A:=\Phi(\rho(B))$, in the latter case we could take $A:=\Phi(\rho(B))\cup D$, where ${\rm reach}\,D>0$, $\dim D=2$ and $\mathrm{dist}(\Phi(\rho(B)),D)>0$.
@@
In the situation of Theorem 1.3 (Theorem L), we can easily observe that the mapping $\Psi:=\Phi\circ\rho|_{B}$ ($\Psi:=\Phi|_{K}$, respectively) is bi-Lipschitz. Using this, we obtain a positive answer to [18, Question 1.7]:
@@
First, given a set $A\subset\mathbb{R}^{d}$ with positive reach and $1\leq k\leq d$, we consider the mapping $\psi_{k}=\psi_{k}^{A}:x\mapsto\operatorname{span}\{\operatorname{Tan}(A,x)\}$ from $T_{k}(A)$ into the Grassmannian $G(d,k)$ and prove its Lipschitzness on some subsets of $T_{k}(A)$ under additional assumptions on $A$. Second, using the Lipschitzness of $\psi_{2}$ on a suitable set, we construct a $C^{1,1}$-diffeomorphism applying Whitney’s $C^{1,1}$ extension theorem to a suitably chosen mapping $f:D\subset\mathbb{R}^{d}\to\mathbb{R}^{d}$. Here our approach differs from that of Lytchak who also uses the Whitney’s $C^{1,1}$ extension, but for a mapping from $\mathbb{R}^{k}$ to $\mathbb{R}^{d-k}$. For our construction of the function $f$ above, we need some basic facts from the theory of product-integration (see [9]), which is the third principal ingredient of our proof.
@@
Moreover, our method also yields the local Lipschitzness of $\psi_{k}$ on $T_{k}(A)$ for any set $A\subset\mathbb{R}^{d}$ with positive reach, which implies that $T_{k}(A)$ lies locally on a $k$-dimensional $C^{1,1}$-surface (Theorem 4.1). (Note that for AA $k$-dimensional, this follows already from Theorem L.) In the case $k=1$ we obtain (using a different method) a stronger version (Theorem 4.4) which will be used substantially in the proof of our main result (Theorem 1.3). Theorem 4.1 easily implies that for $1\leq k<d$, each $k$-dimensional subset of $\mathbb{R}^{d}$ with positive reach can be covered by countably many $k$-dimensional $C^{1,1}$ surfaces (see Corollary 4.2), which is new up to our knowledge.
@@ 2.2. Functions, mappings and surfaces
A mapping $F$ between Banach spaces $X$, $Y$ is called to be $C^{1,1}$ smooth if it has Lipschitz Fréchet derivative on its domain. A mapping $F:U\subset\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a $C^{1,1}$-diffeomorphism if $F$ is injective, both $U$ and $F(U)$ are open and both $F$ and $F^{-1}$ are $C^{1,1}$ smooth.
@@
This result is a rather easy consequence of the classical $C^{1}$ local diffeomorpism theorem and is covered e.g. in [12, Theorem A.9]. It is also an almost immediate consequence of [30, Inverse Function Theorem] (the case $f\in B^{1}_{1}$).
@@
The case $Y=\mathbb{R}$ is an immediate consequence of [15, Corollary 47] (applied with $\omega(t)=t,\ t>0$). For $Y=\mathbb{R}^{n}$ (which is clearly equivalent with the general case) we apply this result coordinate-wisely.
@@
We say that $A\subset\mathbb{R}^{d}$ is a $k$-dimensional $C^{1,1}$ surface ($0<k<d$), if there exist a $k$-dimensional subspace $W$ of $\mathbb{R}^{d}$ and a $C^{1,1}$ mapping $\varphi:W\to W^{\perp}=:V$ such that $A=\{w+\varphi(w):\ w\in W\}$. Then we will also say that $A$ is a $C^{1,1}$ surface associated with $V$.
@@
Slightly reformulating the standard definition (see e.g. [27, Definition 1.30]), we say that a set-valued mapping (or multimapping) $F:x\mapsto F(x)\subset\mathbb{R}^{d}$ defined on a set $C\subset\mathbb{R}^{d}$ is lower semicontinuous (or inner semicontinuous) if for any $x_{0}\in C$, $v_{0}\in F(x_{0})$ and $\varepsilon>0$ there exists $\delta>0$ such that for any $x\in C\cap B(x_{0},\delta)$, $F(x)\cap B(v_{0},\varepsilon)\neq\emptyset$. For convex-valued multimappings, the following result holds (see [26, Theorem 5.9 (a)]):
@@ 2.3. Basic and auxiliary results on sets of positive reach
More precisely, [28, Theorem 15.28, (a), (y)] shows that $A\subset\mathbb{R}^{d}$ is $r$-prox regular if and only if ${\rm reach}\,A\geq r$.
Further, a closed set $A\subset\mathbb{R}^{d}$ is weakly convex w.r.t. $0<r<\infty$ in the sense of [29] (i.e., has Vial property with constant $r$ in the terminology of [28, Definition 16.13]) if and only if $A$ is $r$-prox regular (i.e. ${\rm reach}\,A\geq r$), see [28, Theorem 16.15].
@@
A vector $v\in\mathbb{R}^{d}$ belongs to $\operatorname{Tan}_{C}(A,x)$ (see [26, Definition 6.25], where elements of $\operatorname{Tan}_{C}(A,x)$ are called “regular tangent vectors”; cf. also [27, Theorem 2.2]) if $v$ has the following property:
@@
The normal cone of $A$ at $a$ is defined as the dual cone
@@
For (i) and (ii) see [23, Proposition 3.1]. Assertion (iii) follows e.g. from [23, Proposition 3.1 (vi)] and (6). Statement (iv) was proved by Federer [7], see [23, (3.2), (3.3)]. Assertion (v) follows from [7, Remark 4.15 (1)] (cf. also [23, Lemma 3.4 (ii)]). For (vi), see [29, Proposition 3.8] or [28, Proposition 15.13 (b)] (together with Remark 2.10), and (vii) follows from (vi) using [26, Corollary 6.29]. ∎
@@
Since $a+\operatorname{Tan}(A,a)$ is a closed convex set, there exists a (unique) $c\in P_{a+\operatorname{Tan}(A,a)}(b)$. We have from Proposition 2.12
@@
The intersection of two sets with positive reach need not have positive reach in general, unless additional assumptions are made, as in the following lemma which is a version of [29, Proposition 3.9].
@@
    If $A\cap B$ is compact, ${\rm reach}\,(A,x)>0$ and ${\rm reach}\,(B,x)>0$ for any $x\in A\cap B$ and (11) holds for all $x\in\partial A\cap\partial B$ then ${\rm reach}\,(A\cap B)>0$.
@@
The result follows from [28, Corollary 16.17], Remark 2.10 and the fact that $B$ is $\rho$-strongly convex (see [28, Definition 16.4]), hence also $r$-strongly convex (cf. [28, Corollary 16.12]). ∎
We will use the following well-known fact (see e.g. [28, Proposition 16.20] together with Remark 2.10).
@@
    For every $p,q\in A$ with |pq|<2r|p-q|<2r there exists $s\in A$ such that
    $\left|\frac{p+q}{2}-s\right|\leq r-\sqrt{r^{2}-\frac{|p-q|^{2}}{4}}=:d(p,q,r).$
@@
An elementary computation gives that for |pq|<2r|p-q|<2r,
$\frac{|p-q|^{2}}{8r}\leq d(p,q,r)\leq\frac{|p-q|^{2}}{4r}.$
@@
Assume first that ${\rm reach}\,(A,a)>0$. Using Lemma 2.11 (i) we can find $r>0$ such that ${\rm reach}\,(A,x)>r$ whenever $|x|<r$. Set $\delta:=r/\sqrt{1+K^{2}}$ and $U:=W\cap B(0,\delta)$. Consider any $w\in U$. Then, denoting $x:=(w,f(w))\in\partial A$, we have $|x|<r$ since $f$ is $K$-Lipschitz. By [23, Proposition 3.1 (vi)] there exists $n\in\operatorname{Nor}(A,x)\cap S^{d-1}$ and we have $\overline{B}(x+rn,r)\cap A=\{x\}$ by [22, Lemma 4.3]. Thus the assumptions of [8, Theorem 2.6] are satisfied and we obtain that $f$ is semiconcave on $U$ (recall Remark 2.8).
@@ 3. Lipschitzness of $\psi^{A}_{k}$ on some subsets of $T_{k}(A)$
Our approach (for proving the Lipschitzness of $\psi^{A}_{k}$) is based on the notion of fullness of simplices and on the tangential regularity of sets with positive reach and it is more elementary and also works in more cases than that of [18]. The main idea is contained in the proof of Lemma 3.8. First we present some definitions and several simple lemmas.
@@
for which we know a direct reference only in the complex case ([20, Lemma 3.2]); it is e.g. an obvious consequence of [13, Lemma 2]. Recall also (see [27, Theorems 2.251, 2.250]) that
@@
It follows that $\rho_{k}$ is a metric on $G(d,k)$. It is also well-known that $G(d,k)$ with $\rho_{k}$ is compact (see e.g. [27, Proposition 2.253 and (2.91)]).
@@
Let $x_{0}\in T_{k}(A)$. We can choose $k$ linearly independent vectors $v_{1},\dots,v_{k}\in\operatorname{Tan}(A,x_{0})$; these vectors form a basis of $\widetilde{\operatorname{Tan}}(A,x_{0})$. Now consider an arbitrary $\varepsilon>0$. By [27, Lemma 2.252] there exists $\delta>0$ such that
@@
If $a\in\overline{T_{1}(A)}$ then obviously $a\in A$ and $a\not\in T_{0}(A)$. Further, by (i), $a\not\in T_{k}(A)$ for each $1<k\leq d$, hence $a\in T_{1}(A)$. This proves (iii). Assertion (iv) follows easily from the definition of tangent vectors. ∎
@@
Let $\in T_{1}(A), aba\neq b$. Using Proposition 2.12 we find that there exists $u\in\operatorname{Tan}(A,a)$ such that $|b-a-u|\leq|b-a|^{2}/2r$, hence $\left|\frac{b-a}{|b-a|}-\frac{u}{|b-a|}\right|=\left|\frac{a-b}{|b-a|}-\frac{-u}{|b-a|}\right|\leq\frac{|b-a|}{2r}$. Consequently, using (12) we have $\rho_{1}(\operatorname{span}\{b-a\},\widetilde{\operatorname{Tan}}(A,a))\leq\frac{|b-a|}{2r}.$ Analogously we obtain $\rho_{1}(\operatorname{span}\{a-b\},\widetilde{\operatorname{Tan}}(A,b))\leq\frac{|b-a|}{2r}$, and since $\operatorname{span}\{a-b\}=\operatorname{span}\{b-a\}$, the assertion follows. ∎
@@
Following [31, p. 125] we define the fullness of $\sigma=\sigma(a_{0},\dots,a_{k})$ with ${\rm diam}\,\sigma>0$ as
@@
The first part of (i) is easy and well-known, for the second one see [31, p. 125, (3)]. For (ii) see [31, p. 126, (5)] and (iii) is obvious. The proof of (iv) is easy, cf. [31, p. 126, Lemma 14c] (and the note before it). To prove (v), set $u_{i}:=(a_{i}-a_{0})/|a_{i}-a_{0}|$ and $\lambda_{i}^{*}:=\lambda_{i}|a_{i}-a_{0}|$. Then [31, p. 127, (3)] gives $|\lambda_{i}^{*}|\leq|\sum_{j=1}^{k}\lambda_{j}^{*}u_{j}|/k!\Theta(\sigma)$ and the assertion of (v) follows. ∎
@@
Let $2\leq k\leq d$, $\theta>0$, $v_{1},\dots,v_{k},w\in S^{d-1}$, $\sigma=\sigma(0,v_{1},\dots,v_{k})$ and $\Theta(\sigma)\geq\theta$. Then there exists $1\leq i\leq k$ such that $\Theta(\sigma(0,v_{1},\dots,v_{i-1},w,v_{i+1},\dots,v_{k}))\geq\theta/(k2^{k+1})$.
@@
$\rho_{k}(\widetilde{\operatorname{Tan}}(A,a),V)\leq\frac{L_{k,\theta,r}}{2}|b-a|,$
where $L_{k,\theta,r}:=\frac{k}{(k!\theta)^{2}r}$. By the same way (setting $z_{0}:=b$ and $z_{k}:=a$) we obtain
@@
(19) $\text{the points}\ \ a_{n},b_{n}\ \ \text{are not}\ \ (A,k,2^{-(k+2)}k^{-1}\theta)-\text{related}.$
@@
As an easy consequence we could obtain that $\psi_{k}$ is Lipschitz on any compact subset of $T_{k}(A)$. However, the following lemma (which is important in the proof of our main result) needs a new additional idea.

### v2

@@ 1. Introduction
The following result is a consequence of statements claimed in [7, Remark 4.20] and easily follows from [22, Proposition 7.4].
@@
Recently, A. Lytchak proved an interesting result ([17, Theorem 1.2]) which is essentially an extension of Theorem F. Lytchak’s result deals with sets in $d$-dimensional Riemannian manifold; its (essentially equivalent) Euclidean version is the following.
@@
Note that Theorem L not only generalizes Theorem F (which corresponds to the case $\Phi^{-1}(a)\in{\rm int}\,K$), but also gives (for $k=1$) a local characterization of one-dimensional sets $A$ with positive reach in $\mathbb{R}^{d}$ which is equivalent to that which follows from [22, Corollary 8.9]: on a neighbourhood of $a\in A$, $A$ is either a singleton, or a simple $C^{1,1}$ arc.
If $A\subset\mathbb{R}^{2}$ has positive reach then its local description using semiconcave and semiconvex functions (without any further restrictions) follows easily from [22, Theorem 6.4]. This result, however, cannot be easily extended to higher dimensions.
The main result of the present article is a complete characterization of the local structure of two-dimensional sets $A$ with positive reach in $\mathbb{R}^{d}$ for $d\geq 3$. Note that Theorem L gives such a characterization at points $a\in A$ where $\operatorname{Tan}(A,a)$ has dimension $2$; the only nontrivial remaining case is when $\dim\operatorname{Tan}(A,a)=1$. This case, however, is much different from the case $\dim\operatorname{Tan}(A,a)=2$ - in particular, we cannot expect that $A$ will lie on a two-dimensional $C^{1,1}$ surface in some neighbourhood of $a$. This follows already from [22, Example 7.13], which we will describe now in a slightly generalized version.
@@
Our main result shows (rather surprisingly) that the above example is in a sense universal (up to a $C^{1,1}$-diffeomorphism) for the case when the tangent cone of $A$ at $a$ is a half-line. In order to state our main result (Theorem 1.3 below), we need to include the “both-sided” version of the set $B$ above (where the tangent cone is the full line, see case (ii) in Definition 1.1). (We use the notion $\operatorname{B}$-set as abbreviation of “basic set”.) We also define properly “multirotations” in $\mathbb{R}^{d}$ (Definition 1.2).
@@
Note that if $B$ is a B-set then, since it has positive reach, $T_{1}(B)$ is a compact subset of $\mathbb{R}=\operatorname{span}\{e_{1}\}$ by Lemma 3.3 (iii) (cf. also Remark 5.7 (b)).
@@
On the other hand, if $\Phi$, $B$ and $\rho$ are as above, then $\Phi(\rho(B))$ agrees on a neighbourhood of $a$ with some set $A\subset\mathbb{R}^{d}$ with positive reach satisfying $\dim A\leq 2$ and $\dim(\operatorname{Tan}(A,a))=1$.
@@
It is easy to see that Theorem 1.3 holds if we replace $\dim A\leq 2$” with $\dim A=2$” at both occurrences. The only difference is in the proof of the second part: While in the first case we set in our proof $A:=\Phi(\rho(B))$, in the latter case we could take $A:=\Phi(\rho(B))\cup D$, where ${\rm reach}\,D>0$, $\dim D=2$ and $\mathrm{dist}(\Phi(\rho(B)),D)>0$.
@@
In the situation of Theorem 1.3 (Theorem L), we obtain rather easily that the mapping $\Psi:=\Phi\circ\rho|_{B}$ ($\Psi:=\Phi|_{K}$, respectively) is bi-Lipschitz. Using this, we obtain a positive answer to [17, Question 1.7]:
@@
First, given a set $A\subset\mathbb{R}^{d}$ with positive reach and $1\leq k<d$, we consider the mapping $\psi_{k}=\psi_{k}^{A}:x\mapsto\operatorname{span}\{\operatorname{Tan}(A,x)\}$ from $T_{k}(A)$ into the Grassmannian $G(d,k)$ and prove its Lipschitzness on some subsets of $T_{k}(A)$ under additional assumptions on $A$. Second, using the Lipschitzness of $\psi_{2}$ on a suitable set, we construct a $C^{1,1}$-diffeomorphism applying Whitney’s $C^{1,1}$ extension theorem to a suitably chosen mapping $f:D\subset\mathbb{R}^{d}\to\mathbb{R}^{d}$. Here our approach differs from that of Lytchak who also uses the Whitney’s $C^{1,1}$ extension, but for a mapping from $\mathbb{R}^{k}$ to $\mathbb{R}^{d-k}$. For our construction of the mapping $f$ above, we need some basic facts from the theory of product-integration from [9], which is the third principal ingredient of our proof.
@@
Moreover, our method also yields the local Lipschitzness of $\psi_{k}$ on $T_{k}(A)$ for any set $A\subset\mathbb{R}^{d}$ with positive reach, which implies that $T_{k}(A)$ lies locally on a $k$-dimensional $C^{1,1}$-surface (Theorem 4.1). (Note that for AA $k$-dimensional, this follows already from Theorem L.) In the case $k=1$ we obtain (using a different method) a stronger version (Theorem 4.4) which will be used substantially in the proof of our main result (Theorem 1.3). Theorem 4.1 easily implies that for $1\leq k<d$, each $k$-dimensional subset of $\mathbb{R}^{d}$ with positive reach can be covered by countably many $k$-dimensional $C^{1,1}$ surfaces (see Corollary 4.2), which is a new result up to our knowledge.
@@ 2.2. Functions, mappings and surfaces
A mapping $F$ between Banach spaces $X$, $Y$ is called to be $C^{1,1}$ smooth if it has Lipschitz Fréchet derivative $DF$ on its domain. A mapping $F:U\subset\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a $C^{1,1}$-diffeomorphism if $F$ is injective, both $U$ and $F(U)$ are open and both $F$ and $F^{-1}$ are $C^{1,1}$ smooth.
@@
This result is a rather easy consequence of the classical $C^{1}$ local diffeomorpism theorem and is covered e.g. in [12, Theorem A.9]. It is also an almost immediate consequence of [29, Inverse Function Theorem] (the case $f\in B^{1}_{1}$).
@@
The case $Y=\mathbb{R}$ is an immediate consequence of [15, Corollary 47] (applied with $\omega(t)=t,\ t>0$). For the case $Y=\mathbb{R}^{n}$ (which is clearly equivalent with the general case) we apply this result coordinate-wisely.
@@
Let $0<k<d$. We say that $A\subset\mathbb{R}^{d}$ is a
  1. (i)
    $k$-dimensional $C^{1,1}$ g-surface if there exist a $k$-dimensional subspace $W$ of $\mathbb{R}^{d}$ and a $C^{1,1}$ mapping $\varphi:W\to W^{\perp}$ such that $A=\{w+\varphi(w):\ w\in W\}$,
  2. (ii)
    $k$-dimensional $C^{1,1}$ surface if for any $a\in A$ there exist $\varepsilon>0$ and a $k$-dimensional $C^{1,1}$ g-surface $S\subset\mathbb{R}^{d}$ such that $A\cap B(a,\varepsilon)=S\cap B(a,\varepsilon)$.
Remark 2.6.
$C^{1,1}$ g-surfaces are special types of $C^{1,1}$ surfaces (i.e., $C^{1,1}$ submanifolds of $\mathbb{R}^{d}$) which can be represented as graphs of $C^{1,1}$ mappings. In [22], the term “surface” was used for “g-surface”. All our results involving $C^{1,1}$ surfaces (Theorem 4.1, Corollary 4.2 and Theorem 4.4) could be equivalently formulated with $C^{1,1}$ g-surfaces.
@@
Slightly reformulating the standard definition (see e.g. [26, Definition 1.30]), we say that a set-valued mapping (or multimapping) $F:x\mapsto F(x)\subset\mathbb{R}^{d}$ defined on a set $C\subset\mathbb{R}^{d}$ is lower semicontinuous (or inner semicontinuous) if for any $x_{0}\in C$, $v_{0}\in F(x_{0})$ and $\varepsilon>0$ there exists $\delta>0$ such that for any $x\in C\cap B(x_{0},\delta)$ we have $F(x)\cap B(v_{0},\varepsilon)\neq\emptyset$. For convex-valued multimappings, the following result holds (see [25, Theorem 5.9 (a)]):
@@ 2.3. Basic and auxiliary results on sets of positive reach
More precisely, [27, Theorem 15.28, (a), (y)] shows that $A\subset\mathbb{R}^{d}$ is $r$-prox regular if and only if ${\rm reach}\,A\geq r$.
Further, a closed set $A\subset\mathbb{R}^{d}$ is weakly convex w.r.t. $0<r<\infty$ in the sense of [28] (i.e., has Vial property with constant $r$ in the terminology of [27, Definition 16.13]) if and only if $A$ is $r$-prox regular (i.e. ${\rm reach}\,A\geq r$), see [27, Theorem 16.15].
@@
A vector $v\in\mathbb{R}^{d}$ belongs to $\operatorname{Tan}_{C}(A,x)$ (see [25, Definition 6.25], where elements of $\operatorname{Tan}_{C}(A,x)$ are called “regular tangent vectors”; cf. also [26, Theorem 2.2]) if $v$ has the following property:
@@
The normal cone of $A$ at $a\in A$ is defined as the dual cone
@@
For (i) and (ii) see [22, Proposition 3.1]. Assertion (iii) follows e.g. from [22, Proposition 3.1 (vi)] and (6). Statement (iv) was proved by Federer [7], see [22, (3.2), (3.3)]. Assertion (v) follows from [7, Remark 4.15 (1)] (cf. also [22, Lemma 3.4 (ii)]). For (vi), see [28, Proposition 3.8] or [27, Proposition 15.13 (b)] (together with Remark 2.11), and (vii) follows from (vi) using [25, Corollary 6.29]. ∎
@@
Since $a+\operatorname{Tan}(A,a)$ is a closed set, we can choose $c\in P_{a+\operatorname{Tan}(A,a)}(b)$. We have from Proposition 2.13
@@
The intersection of two sets with positive reach need not have positive reach in general, unless additional assumptions are made, as in the following lemma which is a version of [28, Proposition 3.9].
@@
    If $A\cap B$ is compact, ${\rm reach}\,(A,x)>0$, ${\rm reach}\,(B,x)>0$ for any $x\in A\cap B$ and (11) holds for all $x\in\partial A\cap\partial B$, then ${\rm reach}\,(A\cap B)>0$.
@@
The result follows from [27, Corollary 16.17], Remark 2.11 and the fact that $B$ is $\rho$-strongly convex (see [27, Definition 16.4]), hence also $r$-strongly convex (cf. [27, Corollary 16.12]). ∎
We will use the following well-known fact (see e.g. [27, Proposition 16.20] together with Remark 2.11).
@@
    For every $x,y\in A$ with |xy|<2r|x-y|<2r there exists $s\in A$ such that
    $\left|\frac{x+y}{2}-s\right|\leq r-\sqrt{r^{2}-\frac{|x-y|^{2}}{4}}=:d(x,y,r).$
@@
An elementary computation gives that for |xy|<2r|x-y|<2r,
$\frac{|x-y|^{2}}{8r}\leq d(x,y,r)\leq\frac{|x-y|^{2}}{4r}.$
@@
Assume first that ${\rm reach}\,(A,a)>0$. Using Lemma 2.12 (i) we can find $r>0$ such that ${\rm reach}\,(A,x)>r$ whenever $|x|<r$. Set $\delta:=r/\sqrt{1+K^{2}}$ and $U:=W\cap B(0,\delta)$. Consider any $w\in U$. Then, denoting $x:=(w,f(w))\in\partial A$, we have $|x|<r$ since $f$ is $K$-Lipschitz. By [22, Proposition 3.1 (vi)] there exists $n\in\operatorname{Nor}(A,x)\cap S^{d-1}$ and we have $\overline{B}(x+rn,r)\cap A=\{x\}$ by [21, Lemma 4.3]. Thus the assumptions of [8, Theorem 2.6] are satisfied and we obtain that $f$ is semiconcave on $U$ (recall Remark 2.9).
@@ 3. Lipschitzness of $\psi^{A}_{k}$ on some subsets of $T_{k}(A)$
Our approach (for proving the Lipschitzness of $\psi^{A}_{k}$) is based on the notion of fullness of simplices and on the tangential regularity of sets with positive reach, and it is more elementary and also works in more cases than that of [17]. The main idea is contained in the proof of Lemma 3.8. First we present some definitions and several simple lemmas.
@@
for which we know a direct reference only in the complex case ([19, Lemma 3.2]); however the real case is e.g. an obvious consequence of [13, Lemma 2]. Recall also (see [26, Theorems 2.251, 2.250]) that
@@
It follows that $\rho_{k}$ is a metric on $G(d,k)$. It is also well-known that $G(d,k)$ with $\rho_{k}$ is compact (see e.g. [26, Proposition 2.253 and (2.91)]).
@@
Let $x_{0}\in T_{k}(A)$. We can choose $k$ linearly independent vectors $v_{1},\dots,v_{k}\in\operatorname{Tan}(A,x_{0})$; these vectors form a basis of $\widetilde{\operatorname{Tan}}(A,x_{0})$. Now consider an arbitrary $\varepsilon>0$. By [26, Lemma 2.252] there exists $\delta>0$ such that
@@
Assertion (iv) follows easily from the definition of tangent vectors. If $a\in\overline{T_{1}(A)}$ then obviously $a\in A$ and $a\not\in T_{0}(A)$ by (iv). Further, by (i), $a\not\in T_{k}(A)$ for each $1<k\leq d$, hence $a\in T_{1}(A)$. This proves (iii). ∎
@@
Let $\in T_{1}(A), aba\neq b$. Using Proposition 2.13 we find that there exists $u\in\operatorname{Tan}(A,a)$ such that $|b-a-u|\leq|b-a|^{2}/2r$, hence
$\left|\frac{b-a}{|b-a|}-\frac{u}{|b-a|}\right|=\left|\frac{a-b}{|b-a|}-\frac{-u}{|b-a|}\right|\leq\frac{|b-a|}{2r}.$
Consequently, using (12) we have $\rho_{1}(\operatorname{span}\{b-a\},\widetilde{\operatorname{Tan}}(A,a))\leq\frac{|b-a|}{2r}.$ Analogously we obtain $\rho_{1}(\operatorname{span}\{a-b\},\widetilde{\operatorname{Tan}}(A,b))\leq\frac{|b-a|}{2r}$, and since $\operatorname{span}\{a-b\}=\operatorname{span}\{b-a\}$, the assertion follows. ∎
@@
Following [30, p. 125] we define the fullness of $\sigma=\sigma(a_{0},\dots,a_{k})$ with ${\rm diam}\,\sigma>0$ as
@@
The first part of (i) is easy and well-known, for the second one see [30, p. 125, (3)]. For (ii) see [30, p. 126, (5)] and (iii) is obvious. The proof of (iv) is easy, cf. [30, p. 126, Lemma 14c] (and the note before it). To prove (v), set $u_{i}:=(a_{i}-a_{0})/|a_{i}-a_{0}|$ and $\lambda_{i}^{*}:=\lambda_{i}|a_{i}-a_{0}|$. Then [30, p. 127, (3)] gives $|\lambda_{i}^{*}|\leq|\sum_{j=1}^{k}\lambda_{j}^{*}u_{j}|/k!\Theta(\sigma)$ and the assertion of (v) follows. ∎
@@
Let $2\leq k\leq d$, $\theta>0$, $v_{1},\dots,v_{k},w\in S^{d-1}$, $\sigma=\sigma(0,v_{1},\dots,v_{k})$ and $\Theta(\sigma)\geq\theta$. Then there exists $1\leq i\leq k$ such that
$\Theta(\sigma(0,v_{1},\dots,v_{i-1},w,v_{i+1},\dots,v_{k}))\geq\theta/(k2^{k+1}).$
@@
$\rho_{k}(\widetilde{\operatorname{Tan}}(A,a),V)\leq\frac{L_{k,\theta,r}}{2}|b-a|.$
By the same way (setting $z_{0}:=b$ and $z_{k}:=a$) we obtain
@@
(19) $\text{the points}\ \ a_{n},b_{n}\ \ \text{are not}\ \ (A,k,2^{-(k+2)}k^{-1}\theta)\text{-related}.$
@@
Remark 3.12.
As an easy consequence we obtain using [16, Proposition A48] that $\psi_{k}$ is Lipschitz on any compact subset of $T_{k}(A)$.
However, the following lemma (which is important in the proof of our main result) needs a new additional idea.